# Open-FinLLMs: Open Multimodal Large Language Models for Financial Applications

### v2

@@ Abstract
Financial LLMs hold promise for advancing financial tasks and domain-specific applications. However, they are limited by scarce corpora, weak multimodal capabilities, and narrow evaluations, making them less suited for real-world application. To address this, we introduce Open-FinLLMs, the first open-source multimodal financial LLMs designed to handle diverse tasks across text, tabular, time-series, and chart data, excelling in zero-shot, few-shot, and fine-tuning settings. The suite includes FinLLaMA, pre-trained on a comprehensive 52-billion-token corpus; FinLLaMA-Instruct, fine-tuned with 573K financial instructions; and FinLLaVA, enhanced with 1.43M multimodal tuning pairs for strong cross-modal reasoning. We comprehensively evaluate Open-FinLLMs across 14 financial tasks, 30 datasets, and 4 multimodal tasks in zero-shot, few-shot, and supervised fine-tuning settings, introducing two new multimodal evaluation datasets. Our results show that Open-FinLLMs outperforms afvanced financial and general LLMs such as GPT-4, across financial NLP, decision-making, and multi-modal tasks, highlighting their potential to tackle real-world challenges. To foster innovation and collaboration across academia and industry, we release all codes [^1] https://anonymous.4open.science/r/PIXIU2-0D70/B1D7/LICENSE and models under OSI-approved licenses.
@@
Jimin Huang1, Mengxi Xiao1, Dong Li1, Zihao Jiang1, Yuzhe Yang4, Yifei Zhang5, Lingfei Qian1, Yan Wang1, Xueqing Peng1, Yang Ren1, Ruoyu Xiang1, Zhengyu Chen1, Xiao Zhang1, Yueru He3, Weiguang Han2, Shunian Chen4, Lihang Shen3, Daniel Kim6, Yangyang Yu8, Yupeng Cao8, Zhiyang Deng8, Haohang Li8, Duanyu Feng9, Yongfu Dai9, VijayaSai Somasundaram10, Peng Lu11, Guojun Xiong14, Zhiwei Liu7, Zheheng Luo7, Zhiyuan Yao8, Ruey-Ling Weng1, Meikang Qiu14, Kaleb E Smith15, Honghai Yu5, Yanzhao Lai1, Min Peng2, Jian-Yun Nie11, Jordan W. Suchow8, Xiao-Yang Liu3*, Benyou Wang4*, Alejandro Lopez-Lira10*, Qianqian Xie1,2*, Sophia Ananiadou7,16,17, Junichi Tsujii 1The Fin AI  2Wuhan University  3Columbia University 4The Chinese University of Hong Kong, Shenzhen  5Nanjing University 6Rensselaer Polytechnic Institute  7The University of Manchester 8Stevens Institute of Technology  9Sichuan University  10University of Florida  11University of Montreal 12Yale University  13New York University  14Stony Brook University  15NVIDIA 16Artificial Intelligence Research Centre  17Archimedes/Athena Research Centre xl2427@columbia.edu, wangbenyou@cuhk.edu.cn, alejandro.lopez-lira@warrington.ufl.edu, xqq.sincere@gmail.com * Corresponding Authors
@@ 1 Introduction
The advancements of financial AI have been significantly driven by the progress of natural language processing (NLP) techniques, particularly large language models (LLMs) (Brown et al., 2020; Bubeck et al., 2023). Commercial LLMs like OpenAI’s GPT-4 (Achiam et al., 2023) and open-source LLMs such as Meta AI’s LLaMA series (Touvron et al., 2023a, b) have set new benchmarks in NLP tasks and vertical-domain tasks like medicine, owing to their impressive text understanding and generation capabilities. However, these general-purpose LLMs face limitations in the financial domain due to the knowledge gap. These models are primarily pretrained on general texts and lack an understanding of financial terminology, regulations, and market nuances (Xie et al., 2023a; Wu et al., 2023a; Xie et al., 2023b).
@@
To address current limitations, we introduce Open-FinLLMs, the first open-source multimodal financial LLMs capable of tackling a wide range of financial tasks across text, tabular, time-series, and chart data, excelling in zero-shot, few-shot, and supervised fine-tuning settings. We begin with FinLLaMA, a foundational model pre-trained on a massive 52-billion-token corpus combining text, tabular, and time-series data from high-quality financial sources like reports, papers, and market data. This multi-modal pre-training gives FinLLaMA strong financial insights and analytical skills. We further develop FinLLaMA-Instruct based on FinLLaMA, fine-tuned on 573K diverse, high-quality financial instructions for better instruction-following and broader domain coverage. We also present FinLLaVA, the first model to handle images, text, charts, and tables, trained on 1,430K multimodal financial instruction pairs based on FinLLaMA. Unlike prior models focused on image-text pairs, our fine-tuning incorporates tabular image layouts, improving multi-modal alignment and complex data processing.
@@
In summary, our key contributions are [^2] We will publicly share the training code, datasets, and models.: (1) Introducing Open-FinLLMs, a suite of open-source financial LLMs trained on the largest and most diverse datasets, covering various financial sources and tasks. FinLLaVA is the first open-source model to handle text, tabular, time-series, and chart data. (2) Comprehensive evaluation on 14 financial tasks (30 datasets), 4 multimodal tasks, and three settings (zero-shot, few-shot, and fine-tuning), including two new financial multimodal benchmarks. (3) Extensive results show Open-FinLLMs outperform other financial LLMs, including GPT-4, in financial NLP, decision-making, and multimodal tasks, highlighting their real-world potential.
@@ 2.1 FinLLaMA: Specializing LLaMA3 for Finance with Continual Pre-training
As shown in Table 2, our continual pre-training corpus includes diverse financial data sources for comprehensive coverage: (1) Financial papers: 4 billion tokens from academic papers (2000–2023) on topics like market analysis and economic theory. (2) Conference calls: 5 billion tokens from earnings call transcripts (2004–2021), offering insights into corporate performance. (3) Financial reports: 5 billion tokens from annual and quarterly reports (2005–2020), key for assessing company health. (4) Technical indicators: 12 billion tokens from economic data (2009–2023), essential for investment decisions. (5) News and social media: 7 billion tokens from financial news and social media (1999–2021), providing real-time market trends. (6) Historical data: 13 billion tokens from stock prices and market data (1999–2022), critical for quantitative analysis. (7) SEC filings: 6 billion tokens from SEC reports (1994–2022), detailing corporate activities. Unlike other models, we avoid using web data due to its higher noise level. Further details are in Appendix B.
@@ 2.2 FinLLaMA-Instruct: Domain Task Optimization through Instruction Tuning
Financial Instruction Dataset To optimize FinLLaMA for downstream domain tasks and instruction following ability, we have assembled an extensive instruction-tuning dataset, totaling 573K samples, specifically tailored for financial applications, as shown in Table 3. The dataset is sourced from four open data sources: (1) FLUPE (Xie et al., 2023b), with 123K samples covering key financial NLP tasks; (2) finred (Liu et al., 2023), with approximately 32.67K examples focused on financial report and document comprehension; (3) MathInstruct (Yue et al., 2023), with 262K examples from 13 distinct mathematical rationale datasets; and (4) Sujet-Finance-Instruct-177k, which integrates data from 18 different financial NLP datasets. To ensure the uniqueness and quality of the data, we performed rigorous deduplication and filtering, addressing potential overlaps between datasets like FLUPE and Sujet-Finance-Instruct-177k. See Appendix C for more details on the instruction tuning data and data processing.
@@ 2.3 FinLLaVA: Enabling Multimodal Capabilities via Multimodal Instruction Tuning
Multimodal Instruction Data. We curated a multi-modal dataset combining image, tabular, chart, and text data for comprehensive coverage, as shown in Table 4. For image data, we used ALLaVA-4V (Chen et al., 2024b), LLaVA-v1.5-mix665k (Liu et al., 2024b), and OCR-VQA (Wang et al., 2023). For chart data, we integrated subsets from UniChart (Masry et al., 2023), Chart2Text (Obeid and Hoque, 2020), and ChartQA (Masry et al., 2022). We filtered images relevant to the financial domain using GPT-4o. Unlike prior work (Bhatia et al., 2024), we also included 20k annotated images of tabular data from SynthTabNet (Nassar et al., 2022), and 143k text instructions from Evol-Instruct (Chen et al., 2024b) to improve model generalization and reduce hallucinations. Further details are in Appendix D.2 and D.1.
Multimodal Instruction Finetuning. We use CLIP (Radford et al., 2021) as the visual encoder and FinLLaMA as the language decoder, fine-tuning on our multimodal instruction dataset. Our approach follows LLaVA-1.5’s framework (Liu et al., 2024b), using a two-stage instruction-tuning process: (1) Multimodal Alignment: We align the vision encoder’s output with the language model’s embedding space by freezing both model weights and training a two-layer MLP projector to map visual features into the language model’s space. The training objective maximizes the auto-regressive likelihood: $\sum_{i}^{L}P_{\theta}(X_{i}\mid X_{v},X_{\text{instruct},<i},X_{a,<i})$. (2) Supervised Fine-tuning: In this stage, we update both the language model and MLP projector while keeping the vision encoder frozen, applying the same objective to a new dataset, primarily LLaVA-v1.5-mix665k and Evol-Instruct.
@@ 3.1 Performance of FinLLaMA
In this section, we evaluate the continual pretrained FinLLaMA model using 24 datasets spanning 11 critical financial tasks, categorized into financial-NLP tasks (Sentiment Analysis, Classification, QA, and Named Entity Recognition) and financial-specific tasks (Misinformation, Math, Credit Scoring, Fraud Detection, Financial Distress, Claim Analysis, and Decision Making). We compare its performance against competitor models by reproducing their results where possible or reporting the best publicly available scores.
@@
Zero-shot and Few-shot Performance We evaluate FinLLaMA on both zero-shot and few-shot settings. In the zero-shot setting (Table 8, Appendix F.1), we assess 19 datasets across financial-NLP and financial-specific tasks (Appendix F.4.1, F.4.2). As shown in Table 5 and Figure 5 (Appendix 14), FinLLaMA consistently outperforms baselines, surpassing LLaMA3-8B on all tasks and outperforming BloombergGPT (50B) and LLaMA3.1-8B on most, demonstrating strong zero-shot capabilities from domain-specific continual pre-training. It excels in sentiment analysis, classification, fact-checking (FinFact), and math tasks (MC, KnowledgeMath, DocMath-Eval). In credit scoring, performance varies: on the German dataset, FinLLaMA and LLaMA3-8B predicted one class entirely, while LLaMA3.1-8B predicted the opposite; on the Australian dataset, FinLLaMA performs best; on LendingClub, it beats LLaMA3-8B but not LLaMA3.1-8B. It also shows strong results in fraud detection, financial distress identification, and claim analysis.
In the few-shot setting, we evaluate four datasets across three financial-NLP tasks (Table 9, Appendix F.4.2). As shown in Figure 6, FinLLaMA again outperforms all baselines. In NER, it achieves an F1 of 82.10, surpassing LLaMA3-8B (39.18), LLaMA3.1-8B (49.04), and BloombergGPT (60.82). On FPB, it scores 70.25, outperforming LLaMA3-8B (69.65), LLaMA3.1-8B (13.08), and BloombergGPT (51.07). On FiQA-SA, it achieves 75.34 vs. LLaMA3-8B (52.29), LLaMA3.1-8B (65.39), and BloombergGPT (75.05). On Headlines, it reaches 85.54, outperforming LLaMA3-8B (80.59), LLaMA3.1-8B (59.95), and BloombergGPT (82.20), showing strong classification performance even with minimal examples.
@@
Agent We evaluate FinLLaMA on a financial-specific decision-making task using the FinMem framework (Yu et al., 2024), tested on a single-asset trading dataset from August 15, 2021, to April 25, 2023 (Table 12). This task measures Cumulative Return and Sharpe Ratio as key metrics (Appendix G). As shown in Table 12, FinLLaMA outperforms other LLMs with positive Cumulative Return and Sharpe Ratio, demonstrating profitability in dynamic trading environments. It achieves the highest Sharpe Ratio (above 1), indicating superior risk-return balance. FinLLaMA also maintains stability with an Annual Volatility of 0.4766 and a Maximum Drawdown of 0.2693—both lower than other models. This combination of high Sharpe Ratio and low volatility highlights FinLLaMA’s ability to deliver consistent returns with minimized risk, underscoring the benefits of domain-specific continual pre-training. Detailed results are in Appendix H.

### v3

@@ Abstract
Financial LLMs hold promise for advancing financial tasks and domain-specific applications. However, they are limited by scarce corpora, weak multimodal capabilities, and narrow evaluations, making them less suited for real-world application. To address this, we introduce Open-FinLLMs, the first open-source multimodal financial LLMs designed to handle diverse tasks across text, tabular, time-series, and chart data, excelling in zero-shot, few-shot, and fine-tuning settings. The suite includes FinLLaMA, pre-trained on a comprehensive 52-billion-token corpus; FinLLaMA-Instruct, fine-tuned with 573K financial instructions; and FinLLaVA, enhanced with 1.43M multimodal tuning pairs for strong cross-modal reasoning. We comprehensively evaluate Open-FinLLMs across 14 financial tasks, 30 datasets, and 4 multimodal tasks in zero-shot, few-shot, and supervised fine-tuning settings, introducing two new multimodal evaluation datasets. Our results show that Open-FinLLMs outperforms advanced financial and general LLMs such as GPT-4, across financial NLP, decision-making, and multi-modal tasks, highlighting their potential to tackle real-world challenges. To foster innovation and collaboration across academia and industry, we release all codes and models  [^1] https://huggingface.co/collections/TheFinAI/open-finllms-66b671f2b4958a65e20decbe under OSI-approved licenses.
@@
Jimin Huang1, Mengxi Xiao1, Dong Li1, Zihao Jiang1, Yuzhe Yang4, Yifei Zhang5, Lingfei Qian1, Yan Wang1, Xueqing Peng1, Yang Ren1, Ruoyu Xiang1, Zhengyu Chen1, Xiao Zhang1, Yueru He3, Weiguang Han2, Shunian Chen4, Lihang Shen3, Daniel Kim6, Yangyang Yu8, Yupeng Cao8, Zhiyang Deng8, Haohang Li8, Duanyu Feng9, Yongfu Dai9, VijayaSai Somasundaram10, Peng Lu11, Guojun Xiong14, Zhiwei Liu7, Zheheng Luo7, Zhiyuan Yao8, Ruey-Ling Weng1, Meikang Qiu14, Kaleb E Smith15, Honghai Yu5, Yanzhao Lai1, Min Peng2, Jian-Yun Nie11, Jordan W. Suchow8, Xiao-Yang Liu3*, Benyou Wang4*, Alejandro Lopez-Lira10*, Qianqian Xie1,2*, Sophia Ananiadou7,16,17, Junichi Tsujii 16 1The Fin AI  2Wuhan University  3Columbia University 4The Chinese University of Hong Kong, Shenzhen  5Nanjing University 6Rensselaer Polytechnic Institute  7The University of Manchester 8Stevens Institute of Technology  9National University of Singapore 10University of Florida  11University of Montreal  12Yale University  13New York University 14Harvard University  15NVIDIA 16Artificial Intelligence Research Centre  17Archimedes/Athena Research Centre xl2427@columbia.edu, wangbenyou@cuhk.edu.cn, alejandro.lopez-lira@warrington.ufl.edu, xqq.sincere@gmail.com * Corresponding Authors
@@ 1 Introduction
The advancements of financial AI have been significantly driven by the progress of natural language processing (NLP) techniques, particularly large language models (LLMs) (Brown et al., 2020; Bubeck et al., 2023). Commercial LLMs like OpenAI’s GPT-4 (Achiam et al., 2023) and open-source LLMs such as Meta AI’s LLaMA series (Touvron et al., 2023a, b) have set new benchmarks in NLP tasks and vertical-domain tasks like medicine, owing to their impressive text understanding and generation capabilities. However, these general-purpose LLMs face limitations in the financial domain due to the knowledge gap. These models are primarily pretrained on general texts and lack an understanding of financial terminology, regulations, and market nuances (Xie et al., 2023a; Wu et al., 2023a; Xie et al., 2023b). Additionally, they are unable to effectively handle non-text data, such as tabular and time-series data, which are critical components of financial knowledge.
@@
To address these limitations, in this paper, we introduce Open-FinLLMs, a series of financial large language models tailored for various financial tasks. We begin with FinLLaMA, a groundbreaking foundational model pre-trained on a massive 52-billion-token corpus comprising text, tabular, and time-series data from high-quality financial sources such as reports, papers, and market data for the first time. This innovative pre-training strategy incorporating extensive data modalities equips FinLLaMA with deep financial insights and analytical capabilities. We further develop FinLLaMA-Instruct by fine-tuning the model with expanded datasets of 573K diverse and high-quality financial instructions. Enriched instruction-tuning datasets enhance not only the model’s ability to follow instructions but also coverage of financial domain knowledge, leading to better performance for a wide range of downstream tasks. To handle multimodal financial data, we present FinLLaVA, leveraging 1,430K financial multimodal instruction pairs, including images, text, charts, and tabular data. Unlike traditional methods that primarily focus on standard image-text pairs, our fine-tuning process is the first to incorporate chart image-text pairs and images of tabular layouts. This innovation enables the model to effectively interpret and process complex financial data with improved precision and versatility.
@@
In summary, our key contributions are [^2] We will publicly share the training code, datasets, and models.: (1) We introduce Open-FinLLMs, a groundbreaking suite of financial LLMs trained on three comprehensive datasets tailored for different training stages. This structured training pipeline ensures a robust understanding of financial terminology, numerical data, and complex financial contexts. (2) By pioneering the integration of tabular and time-series data, Open-FinLLMs are the first financial foundation models with advanced multimodal capabilities. This multimodal approach bridges the gap between textual and structured financial data, enabling a more comprehensive capture of financial knowledge from diverse data types. (3) Extensive experiments demonstrate Open-FinLLMs’ superior performance across a range of financial and multimodal benchmark tasks, including zero-shot and few-shot settings. Notably, the models excel in both traditional NLP tasks and finicial tasks in real applications, showcasing their readiness to tackle complex challenges understanding multimodal finicial data. (4) We have open-sourced the training code, datasets, and models, providing a valuable reference for future research on LLMs in the financial domain. The shared code and training methodologies can also support advancements in other fields that require processing time-series data or understanding multimodal information.
2 Related Work
2.1 Financial Large Language Models
2.2 Domain Specialization of Large Language Models
@@ 3.1 FinLLaMA: Specializing LLaMA3 for Finance with Continual Pre-training
As shown in Table 2, our continual pre-training corpus encompasses a wide array of data sources to ensure comprehensive coverage of financial knowledge: (1) Financial papers: $4$ billion tokens from academic papers and research articles, offering a strong foundation in financial concepts and theories. These papers, sourced from SSRN [^3] https://www.ssrn.com/index.cfm/en/ and open-source conference proceedings, span from 2000 to 2023 and cover topics such as market analysis, financial modeling, and economic theory. (2) Conference calls: 5 billion tokens of open-source transcripts from earnings calls, analyst meetings, and investor briefings, collected from 09/08/2004 to 12/17/2021, providing real-time insights into corporate performance and strategies. (3) Financial reports: This component consists of 5 billion tokens from annual and quarterly reports, covering the period from 2005 to 2020, crucial for assessing a company’s financial health and market positioning. (4) Technical indicators: 12 billion tokens of economic indicators and financial ratios sourced from company filings and open market data spanning from 2009 to 2023, essential for macroeconomic analysis and investment decisions. (5) News and social media: 7 billion tokens from financial news outlets and social media platforms collected from 1999 to 2021, offering timely updates on market trends and public opinion. (6) Historical data: 13 billion tokens of historical stock prices, trading volumes, and market data from 1999 to 2022, vital for quantitative analysis and algorithmic trading. (7) SEC filings: 6 billion tokens from U.S. SEC filings, such as 10-K and 10-Q reports spanning from 1994 to 202, providing comprehensive insights into corporate activities and performance. Unlike existing models, we chose not to use web data due to its higher noise level compared to other data sources. See Appendix A for more details of the pre-training data.
@@ 3.2 FinLLaMA-Instruct: Domain Task Optimization through Instruction Tuning
Financial Instruction Dataset To optimize FinLLaMA for downstream domain tasks and instruction following ability, we have assembled an extensive instruction-tuning dataset, totaling 573K samples, specifically tailored for financial applications, as shown in Table 3. The dataset is sourced from four open data sources: (1) FLUPE (Xie et al., 2023b), with 123K samples covering key financial NLP tasks; (2) finred (Liu et al., 2023), with approximately 32.67K examples focused on financial report and document comprehension; (3) MathInstruct (Yue et al., 2023), with 262K examples from 13 distinct mathematical rationale datasets; and (4) Sujet-Finance-Instruct-177k, which integrates data from 18 different financial NLP datasets. To ensure the uniqueness and quality of the data, we performed rigorous deduplication and filtering, addressing potential overlaps between datasets like FLUPE and Sujet-Finance-Instruct-177k. See Appendix B for more details on the instruction tuning data and data processing.
@@ 3.3 FinLLaVA: Enabling Multimodal Capabilities via Multimodal Instruction Tuning
Multimodal Instruction Data. We curated a diverse multi-modal dataset comprising image, tabular, chart, and text data to ensure comprehensive coverage of various data formats, as shown in Table 4. For image data, we utilized three vision instruction datasets: ALLaVA-4V (Chen et al., 2024b), LLaVA-v1.5-mix665k (Liu et al., 2024b), and OCR-VQA (Wang et al., 2023). For chart data, we integrate subsets from multiple sources: UniChart (Masry et al., 2023), with 5k chart image-text pairs; Chart2Text (Obeid and Hoque, 2020), with 30k chart image-text pairs; and ChartQA (Masry et al., 2022), featuring 20k chart images and their associated question-answering (QA) pairs. We used GPT-4o to evaluate these datasets and filtered out images most relevant to the financial domain. Further details are provided in Appendix C.2. Different from previous work (Bhatia et al., 2024), which only image and chart data, we selected a subset of the SynthTabNet (Nassar et al., 2022) dataset, consisting of 20k annotated images of data in tabular layouts. Further details are provided in Appendix C.1. Additionally, we included Evol-Instruct (Chen et al., 2024b), a dataset of 143k pure text instructions, to enhance the model’s generalization capabilities and reduce the risk of hallucination.
Multimodal Instruction Finetuning. We utilize CLIP (Radford et al., 2021) as our visual encoder in conjunction with the FinLlaMA language decoder, fine-tuning the model on our multimodal instruction dataset. Our approach follows the training framework established by LLaVA-1.5 (Liu et al., 2024b), implementing a two-stage instruction-tuning process. Multimodal Alignment: In this initial stage, we aim to align the vision encoder’s output with the language model’s embedding space. During this phase, both the vision encoder and LLM weights remain frozen. The key objective is to train a two-layer MLP projector to bridge the gap between the vision encoder’s features and the LLM’s embedding. For each input, consisting of an image $X_{v}$, instructions $X_{\text{instruct}}$ that may involve single-turn or multi-turn conversations, and the target answer $X_{a}$, the vision encoder processes the image data to generate a vision feature: $Z_{v}=g(X_{v})$. The MLP projector then maps $Z_{v}$ into the embedding space of the language model: $H_{v}=f_{\text{MLP}}(Z_{v};\theta)$, where $\theta$ represents the trainable parameters of the projector. The training objective is to maximize the auto-regressive likelihood: $\sum_{i}^{L}P_{\theta}(X_{i}\mid X_{v},X_{\text{instruct},<i},X_{a,<i})$, where $L$ is the sequence length of the target answer $X_{a}$, and $X_{\text{instruct},<i}$ and $X_{a,<i}$ are the tokens of instructions and answers preceding the current prediction $X_{i}$. Supervised Fine-tuning: In the second stage, we continue updating the parameters of both the language model and the MLP projector, while keeping the vision encoder’s parameters frozen. We maintain the same autoregressive training objective as the previous stage but apply it to a different dataset. As shown in Table 4, we utilize LLaVA-v1.5-mix665k and Evol-Instruct as our primary training data sources.
@@ 4.1 Performance of FinLLaMA
In this section, we evaluate the continual pretrained FinLLaMA model using 26 datasets spanning 11 critical financial tasks, categorized into financial-NLP tasks (Sentiment Analysis, Classification, QA, and Named Entity Recognition) and financial-specific tasks (Misinformation, Math, Credit Scoring, Fraud Detection, Financial Distress, Claim Analysis, Decision Making, Auditing and Risk & Compliance). We compare its performance against competitor models by reproducing their results where possible or reporting the best publicly available scores.
@@
Zero-shot Performance In zero-shot scenarios, as shown in Table 8, our evaluation uses 19 datasets covering both financial-NLP and financial-specific tasks, A detailed description of evaluation tasks and detailed prompts for each dataset are available in Appendix E.4.1 and Appendix E.4.2.
Overall, Table 5 (and Figure 5 in Appendix 14) demonstrates that FinLLaMA outperformed the baseline models on most financial tasks, highlighting its robustness and versatility in zero-shot settings. It surpasses its backbone model, LLaMA3-8B, on all tasks, highlighting the effectiveness of continual pre-training with large-scale domain-specific data in enhancing financial knowledge. Additionally, FinLLaMA exceeds the performance of BloombergGPT, despite its larger model size of 50B, and also outperforms the current most capable open-source LLM, LLaMA3.1-8B on most tasks.
FinLLaMA shows exceptional performance in sentiment analysis and classification tasks, demonstrating its proficiency in fundamental financial operations. Its improved accuracy in fact-checking, as shown by the FinFact dataset, highlights its ability to comprehend and evaluate financial information, enabling precise judgments on claims. In math problem-solving tasks, our model shows improvement across all datasets (MC, KnowledgeMath, and DocMath-Eval) compared to LLaMA3-8B.
In credit scoring tasks, performance varies significantly across datasets. Upon manual inspection of prediction results, we found that on the German dataset, both our model and LLaMA3-8B predicted all cases as one category, while LLaMA3.1-8B predicted the opposite. On the Australian dataset, our model’s superior performance demonstrates the benefits of continuous pre-training, even with anonymized features. In the LendingClub dataset, our model outperformed LLaMA3-8B, though not LLaMA3.1-8B, likely due to its larger scale training data. Additionally, FinLLaMA excelled in fraud detection, financial distress identification, and claim analysis, showcasing its robust capabilities across diverse financial tasks.
Few-shot Performance For the few-shot evaluation setting, we use four datasets covering three financial-NLP tasks, as shown in Table 9. These tasks are aligned with the BloombergGPT evaluation settings to ensure consistency and comparability. Detailed descriptions of the evaluation tasks can be found in Appendix E.4.2.
Overall, as shown in Fig. 6, FinLLaMA consistently outperforms baseline models across a wide range of financial tasks, demonstrating its robustness and versatility in few-shot settings. In the NER task, FinLLaMA achieves a remarkable F1 score of 82.10, significantly surpassing its backbone model LLaMA3-8B (39.18), LLaMA3.1-8B (49.04), and BloombergGPT (60.82). This highlights the substantial improvement in entity recognition due to continual pre-training. For the FPB dataset, FinLLaMA achieves an F1 score of 70.25 in few-shot settings, outperforming LLaMA3-8B (69.65), LLaMA3.1-8B (13.08), and BloombergGPT (51.07). Similarly, in the FiQA-SA dataset, FinLLaMA scored 75.34, surpassing LLaMA3-8B (52.29), LLaMA3.1-8B (65.39), and BloombergGPT (75.05). In the Headlines dataset for classification, FinLLaMA achieves a score of 85.54, outperforming LLaMA3-8B (80.59), LLaMA3.1-8B (59.95), and BloombergGPT (82.20). These results demonstrate FinLLaMA’s strong ability to classify financial texts accurately with minimal examples.
@@
Agent We further assess the performance of FinLLaMA on a financial-specific task (Decision-Making) using the FinMem agent framework (Yu et al., 2024), tested on a single-asset trading dataset comprising multi-sourced financial data from August 15, 2021, to April 25, 2023. The results are presented in Table 12. This task assesses the LLM’s proficiency in single-asset trading, with Cumulative Return and Sharpe Ratio as the key performance metrics. For more tasks and datasets, please refer to Appendix F. As shown in Table 12, FinLLaMA outperforms other LLMs with positive Cumulative Return and Sharpe Ratio metrics, demonstrating profitability in dynamic trading environments. It achieves the highest Sharpe Ratio (SR) of over 1, indicating a superior risk-return balance. Additionally, FinLLaMA maintains investment stability with an Annual Volatility of 0.4766 and a Maximum Drawdown of 0.2693, both lower than other models. This combination of high Sharpe Ratio and low volatility highlights FinLLaMA’s ability to deliver consistent returns with minimized risk, making it highly reliable for trading strategies. These results highlight the significant impact of continual pre-training in enhancing FinLLaMA’s performance. For detailed results, please refer to Appendix G.